OpenStudy (darkigloo):

if the substitution u=x/2 is made, the integral ∫[2,4] ( 1 - (x/2)^2 ) / x dx = ?

OpenStudy (anonymous):

\[\int_2^4\frac{1-(\frac{x}{2})^2}{x}dx\]?

OpenStudy (darkigloo):

yes

OpenStudy (anonymous):

isn't this the same as \[\int_2^4(\frac{1}{x}-\frac{x}{4})dx\]?

OpenStudy (darkigloo):

so what happened to the x in x^2/4?

OpenStudy (anonymous):

cancelled with the \(x\) in the denominator

OpenStudy (darkigloo):

ohh ok i see.

OpenStudy (anonymous):

if i am reading it right, it is \[\frac{1-\left(\frac{x}{2}\right)^2}{x}=\frac{1-\frac{x^2}{4}}{x}=\frac{1}{x}-\frac{x}{4}\]

OpenStudy (darkigloo):

right.

OpenStudy (anonymous):

k so no substitution for this one

OpenStudy (darkigloo):

\[\int\limits_{1}^{2} \frac{ 1-u^2 }{ u } du\] \[\int\limits_{2}^{4} \frac{ 1-u^2 }{ u } du\] \[\int\limits_{1}^{2} \frac{ 1-u^2 }{ 2u } du\] \[\int\limits_{1}^{2} \frac{ 1-u^2 }{ 4u } du\] \[\int\limits_{2}^{4} \frac{ 1-u^2 }{ 2u } du\]

OpenStudy (darkigloo):

those are the answer choices. i don't know what to do..

OpenStudy (anonymous):

is see

OpenStudy (anonymous):

if you put \(u=\frac{x}{2}\) then \[2du=dx\] and \(x=2u\) also \(u(2)=1, u(4)=2\)

OpenStudy (anonymous):

this gives \[\int \frac{1-u^2}{2u}2du\] and the 2's cancel leaving you with \[\int_1^2\frac{1-u^2}{u}du\]

OpenStudy (darkigloo):

Thank you!